# Groupoid Graded Semisimple Rings

### v1

@@ 5.5. On the uniqueness of the representation as matrix rings
and it follows that $\pi$ is surjective too. ∎
@@
Let $H$ be a $\Gamma$-graded ring and $\overline{\Sigma}:=(\Sigma_{i})_{i\in I}\in\mathcal{P}(\Gamma)^{I}$ be a fully matricial sequence for $H$. Let $R:=\operatorname{M}_{I}(H)(\overline{\Sigma})$. Fix $i_{0}\in I$ and consider $M:=\bigoplus\limits_{\sigma\in\Sigma_{i_{0}}}(E_{i_{0}i_{0}}^{r(\sigma)}R)(\sigma^{-1})$. The following assertions hold true.
@@
When $\Gamma$ is a group, Theorem 5.45 can be stated as follows: Let $G$ be a group, $D,D^{\prime}$ be $G$-graded division rings and $\overline{\sigma}:=(\sigma_{1},...,\sigma_{n})\in G^{n}$, $\overline{\delta}:=(\delta_{1},...,\delta_{m})\in G^{m}$. Then $\operatorname{M}_{n}(D)(\overline{\sigma})\cong_{gr}\operatorname{M}_{m}(D^{\prime})(\overline{\delta})$ if and only if $n=m$ and there exist a permutation $\pi$ of $\{1,...,n\}$ and $\tau\in G$ such that $D^{\prime}\cong_{gr}\operatorname{M}_{1}(D)(\tau^{-1})$ and $\delta_{\pi(i)}\in\tau(\operatorname{supp}D)\sigma_{i}$ for each $i=1,...,n$. This is essentially what was achieved in [11, p. 32-33]. Therefore, Theorem 5.45 generalizes the group graded case and we believe that our proof here, when adapted to $\Gamma$ being a group, provides a more elementary proof (or at least based on more elementary facts) of the result of [11]. ∎
@@ 6. The graded Jacobson-Chevalley density theorem
We begin extending the definitions of gr-simple rings and modules for not necessarily object unital rings. Let $R$ be a $\Gamma$-graded ring and $M$ be a $\Gamma$-graded $R$-module. We say that $M$ is gr-simple if $MR\neq 0$ and the only graded submodules of $M$ are $\{0\}$ and $M$. And $M$ is said to be faithful if its right annihilator is zero, that is,
@@
Let $D$ be a $\Gamma$-graded division ring and $V$ be a $\Gamma$-graded unital left $D$-module. A graded subring $T$ of $\operatorname{END}_{D}(V)$ is said to be a gr-dense subring of $\operatorname{END}_{D}(V)$ if for all $n>0$, $\gamma_{1},...,\gamma_{n}\in\Gamma$, pseudo-linearly independent sequence $(u_{1},...,u_{n})\in V_{\gamma_{1}}\times\cdots\times V_{\gamma_{n}}$ and $(v_{1},...,v_{n})\in V_{\delta_{1}}\times\cdots\times V_{\delta_{n}}$ with $r(\delta_{i})=r(\gamma_{i})$ for all $i=1,...,n$ there exists $t\in T$ such that $(u_{i})t=v_{i}$ for all $i=1,...,n$. By Corollary 4.5, $\operatorname{END}_{D}(V)$ is a gr-dense subring of $\operatorname{END}_{D}(V)$. In fact, if $(u_{1},...,u_{n})\in V_{\gamma_{1}}\times\cdots\times V_{\gamma_{n}}$ is a pseudo-linearly independent sequence and $(v_{1},...,v_{n})\in V_{\delta_{1}}\times\cdots\times V_{\delta_{n}}$ with $r(\delta_{i})=r(\gamma_{i})$ for all $i=1,...,n$, then, for each $i=1,...,n$, there exists $t_{i}\in\operatorname{END}_{D}(V)_{\gamma_{i}^{-1}\delta_{i}}$ such that $(u_{i})t_{i}=v_{i}$ and $(u_{j})t_{i}=0$ for every $j\neq i$. It suffices to take $t:=t_{1}+\cdots+t_{n}$.
@@
Now suppose that $n\geq 1$ and that the result holds for $n-1$. Let ${}_{D}V$ be a graded $D$-submodule of ${}_{D}S$ with $\operatorname{pdim}_{D}V=n$ and $x\in\operatorname{h}(S)\setminus V$. Let $(v_{1},...,v_{n})$ be a pseudo-basis of ${}_{D}V$. If $n>1$, define
@@
for all $a\in A$. Then the following homomorphism of $R$-modules is well defined
@@
$(4)\implies(1):$ It follows from Theorem 5.23. ∎
@@ 7. Pseudo free module rings
    Conversely, suppose that $R$ is a gr-semisimple ring. Then $R$ is a pfm ring if, and only if, every gr-simple right $R$-module is pseudo-free.
@@
{graded division rings}{pfm rings}{gr-semisimple $\left\{\begin{array}[]{c}\textrm{graded division rings}\end{array}\right\}\subset\left\{\begin{array}[]{c}\textrm{pfm rings}\end{array}\right\}\subset\left\{\begin{array}[]{c}\textrm{gr-semisimple rings}\end{array}\right\}.$
@@
We point out that, in Example 7.2(2), $\mathcal{B}_{1}=\{\mathbb{I}_{(1,1)}\}$ and $\mathcal{B}_{2}=\{E_{13},\,E_{23}\}$ are two pseudo-bases of the $\Gamma$-graded $R$-module $S=R((1,1))$. Thus, we don’t have uniqueness of the cardinality of pseudo-bases of $\Gamma$-graded $R$-modules over pfm rings, and as a consequence, over gr-semisimple rings. Such uniqueness will characterize gr-division rings, but we will also be able to define an invariant similar to dimension for pfm rings, see Section 7.2.
@@ 7.1. Characterization of pfm rings
Fix $j\in J$ and $k\in K_{j}$. Set $R_{j}:=\operatorname{M}_{K_{j}}(D_{j})(\overline{\sigma}_{j})$ and let $\iota_{j}:R_{j}\to R$ be the canonical inclusion. Since the $\Gamma$-graded right $R$-module $\iota_{j}(E_{kk}^{e_{j}}R_{j})$ has a pseudo-basis, there exist $\gamma_{j}\in\Gamma$ and $a_{j}\in(E_{kk}^{e_{j}}R_{j})_{\gamma_{j}}$ such that $\iota_{j}(a_{j})\cdot r\neq 0$ for all $r\in 1_{d(\gamma_{j})}R\setminus\{0\}$. In particular, $a_{j}\neq 0$ and, thus, $K_{j,d(\gamma_{j})}\neq\emptyset$. Let $j^{\prime}\in J$ be such that there exists $k^{\prime}\in K_{j^{\prime},d(\gamma_{j})}$. Then $0\neq E_{k^{\prime}k^{\prime}}^{e_{j^{\prime}}}\in(R_{j^{\prime}})_{d(\gamma_{j})}$ and it follows that $\iota_{j}(a_{j})\iota_{j^{\prime}}(E_{k^{\prime}k^{\prime}}^{e_{j^{\prime}}})\neq 0$. But this is only possivel if $j=j^{\prime}$. Hence, $\{j^{\prime}\in J:K_{j^{\prime},f_{j}}\neq\emptyset\}=\{j\}$. Now observe that $a_{j}\cdot r_{j}\neq 0$ for all $r_{j}\in R_{j}(d(\gamma_{j}))\setminus\{0\}$. Proceeding as in the proof of $(2)\implies(3)$ of Theorem 7.3, we get that $K_{j,d(\gamma_{j})}$ contains exactly one element.
@@
$(6)\implies(2)$: Let $j\in J$. By Theorem 5.23, there exist a nonempty set $K_{j}$, $e_{j}\in\Gamma_{0}$, a $\Gamma$-graded division ring $D_{j}$ with $\operatorname{supp}(D_{j})\subseteq e_{j}\Gamma e_{j}$ and a $d$-finite sequence $\overline{\sigma}_{j}:=(\sigma_{jk})_{k\in K_{j}}\in(e_{j}\Gamma)^{K_{j}}$ such that
@@ 7.2. Invariance of the number of elements of pseudo-bases
As we are going to see next, it is still possible to define an invariant for graded modules over pfm rings similar to the pseudo-dimension and that coincides with the graded length of a finitely generated graded module over a pfm ring. For that we begin pointing out some facts about gr-semisimple modules.
@@
Suppose now that $X$ is a $\Gamma$-graded module. We say that a pseudo-linearly independent sequence $(x_{i})_{i\in I}$ of homogeneous elements of $X$ is a gr-simple sequence if $x_{i}R$ is a gr-simple $R$-module for all $i\in I$. If, moreover, $(x_{i})_{i\in I}$ is a pseudo-basis of $X$ we say that it is a gr-simple pseudo-basis of $X$. In this event, we will write $\operatorname{spdim}_{R}(X)=|I|$. We proceed to show that $\operatorname{spdim}_{R}(X)$ is a well behaved invariant for pfm rings.
@@
Let $R=\bigoplus\limits_{\gamma\in\Gamma}R_{\gamma}$ be a pfm ring and $M=\bigoplus\limits_{\gamma\in\Gamma}M_{\gamma}$ be a $\Gamma$-graded $R$-module. The following assertions hold true.
@@
    If $N$ is a graded submodule of $M$, then
    $\operatorname{spdim}_{R}(N)+\operatorname{spdim}_{R}(M/N)=\operatorname{spdim}_{R}(M).$
@@ 7.3. More on gr-division rings
    There exist a set $J$, a family of nonempty subsets $\{K_{j}:j\in J\}$, a sequence of idempotents $(e_{j})_{j\in J}\in(\Gamma_{0})^{J}$ such that
@@
$(3)\implies(4)$: Clear.
$(4)\implies(5)$: By Theorem 5.30, there exist families $\{K_{j}:j\in J\}$ and $\{e_{j}:j\in J\}$ of nonempty sets and idempotents of $\Gamma$, respectively, and there exist, for each $j\in J$, a $\Gamma$-graded division ring $D_{j}$ with $\operatorname{supp}(D_{j})\subseteq e_{j}\Gamma e_{j}$ and a $d$-finite sequence $\overline{\sigma}_{j}:=(\sigma_{jk})_{k\in K_{j}}\in(e_{j}\Gamma)^{K_{j}}$, such that the family $\{\operatorname{M}_{K_{j}}(D_{j})(\overline{\sigma}_{j}):j\in J\}$ is summable and
@@
$(5)\implies(6)$: For each $j\in J$, $R_{j}:=\operatorname{M}_{K_{j}}(D_{j})(\overline{\sigma}_{j})$ is a gr-prime gr-division ring, by Theorem 4.2. Suppose now that $j,j^{\prime}\in J$ and $\gamma\in\operatorname{supp}(R_{j})\cap\operatorname{supp}(R_{j^{\prime}})$. Since $|K_{j,e}|=1$ for all $e\in\Gamma^{\prime}_{0}(R)$, there exists a unique $p\in K_{j}$ such that $d(\sigma_{jp})=r(\gamma)$. Analogously, there exists a unique $p^{\prime}\in K_{j^{\prime}}$ such that $d(\sigma_{j^{\prime}p^{\prime}})=r(\gamma)$. Thus, we obtain nonempty $K_{j,r(\gamma)}$ and $K_{j^{\prime},r(\gamma)}$, hence $j,j^{\prime}\in J_{r(\gamma)}$ and, therefore, $j^{\prime}=j$.
@@
The next result says that to be a pfm ring is equivalent to be a gr-division ring for object crossed products.
@@
The following assertions hold true.
@@ 8. Semisimple categories
Many of the results in this section can be proved in a more general context. The authors are working in a paper where they introduce the concept of groupoid graded categories, which generalize group graded categories, and results are obtained for such more general categories.
@@
In [29], $\mathcal{C}(A,-)$ and $\mathcal{C}(-,A)$ are denoted by $\mathcal{C}_{A}$ and $\mathcal{C}_{A}^{~{}*}$, respectively. In [29, p. 18], a right (left) ideal of $\mathcal{C}$ is defined as a subfunctor of $\mathcal{C}_{A}$ (of $\mathcal{C}_{A}^{~{}*}$) for some $A\in\mathcal{C}_{0}$. We point out that we are using the opposite notation because our composition of morphisms in the category $\mathcal{C}$ is the usual from right to left, while [29, p. 7] composes morphisms of $\mathcal{C}$ from left to right.
In what follows, inspired by [15, Proposition 2 (p. 347)], we present a possible motivation for the previous definitions.
@@
If $M=M(\varepsilon_{A})$ for some $A\in\mathcal{C}_{0}$, then we will write $M(-,A)$ instead of $M(-,\mathcal{C}_{0})$.
@@
The constructions of the previous paragraphs are related by the next result. Recall that, by Remark 2.5(1), every unital right $R[\mathcal{C}]$-module $M$ can be regarded as a $\mathcal{G}$-graded $R[\mathcal{C}]$-module making $M=M(\varepsilon_{A})$ for some $\varepsilon_{A}\in\mathcal{G}_{0}$. Therefore, Theorem 8.1(1) is a rephrasing of [15, Proposition 2 (p. 347)].
@@
Let $\mathcal{G}:=\mathcal{C}_{0}\times\mathcal{C}_{0}$. There exists an additive covariant isomorphism of categories
$\mathcal{G}-\operatorname{gr\textrm{-}\!}R[\mathcal{C}]$ $\longrightarrow\operatorname{Fun_{gr}}(\mathcal{C}^{op},\mathcal{G}\varepsilon-\mathbb{Z}^{(\mathcal{G}_{0})}\operatorname{\!\textrm{-}mod})$
$M$ $\longmapsto M(-,\mathcal{C}_{0})$
whose inverse is given by $F\longmapsto M[F]$. If $A\in\mathcal{C}_{0}$ , then such isomorphism restricts to
@@
    an additive covariant isomorphism of categories
    $\varepsilon_{A}\mathcal{G}-\operatorname{gr\textrm{-}\!}R[\mathcal{C}]$ $\longrightarrow\operatorname{Fun}(\mathcal{C}^{op},\mathcal{A}b)$
    $M$ $\longmapsto M(-,A).$
@@
    that sends $R[\mathcal{C}](\varepsilon_{A})$ to $\mathcal{C}(-,A)$.
@@
[…]
@@
Observe that, for each $\varepsilon\in\mathcal{G}_{0}$, the correspondences
$\mathcal{G}-\operatorname{gr\textrm{-}\!}R[\mathcal{C}]$ $\longrightarrow\operatorname{Fun_{gr}}(\mathcal{C}^{op},\mathcal{G}\varepsilon-\mathbb{Z}^{(\mathcal{G}_{0})}\operatorname{\!\textrm{-}mod})$
$M$ $\longmapsto M(-,\mathcal{C}_{0})$
$M[F]$ $\longmapsfrom F$
are one inverse of the other.
(1) Fix $A\in\mathcal{C}_{0}.$ For $F\in\operatorname{Fun}(\mathcal{C}^{op},\mathcal{A}b)$, consider $M[F]$ as an object of $\varepsilon_{A}\mathcal{G}-\operatorname{gr\textrm{-}\!}R[\mathcal{C}]$ via $M[F]_{(A,X)}:=F(X)$ for each $X\in\mathcal{C}_{0}$. Then $F\longmapsto M[F]$ defines the inverse of the functor $M\longmapsto M(-,A)$.
(2) Note that if $M$ is a graded right ideal of $R[\mathcal{C}]$ with $M=M(\varepsilon_{A})$ for some $A\in\mathcal{C}_{0}$, then $M(-,A)$ is a subfunctor of $\mathcal{C}(-,A)$, i.e., a right sieve on $A$ in $\mathcal{C}$. And, if $F$ is a right sieve on $A$ in $\mathcal{C}$, then $M[F]$ is a graded right ideal of $R[\mathcal{C}]$ contained in $R[\mathcal{C}](\varepsilon_{A})$. Thus, (2) follows from (1). ∎

### v2

@@ 5.5. On the uniqueness of the representation as matrix rings
and it follows that $\pi$ is surjective as well. ∎
@@
Let $H$ be a $\Gamma$-graded ring and $\overline{\Sigma}:=(\Sigma_{i})_{i\in I}\in\mathcal{P}(\Gamma)^{I}$ be a fully matricial sequence for $H$. Let $R:=\operatorname{M}_{I}(H)(\overline{\Sigma})$. Fix $i_{0}\in I$ and consider $M:=\bigoplus\limits_{\sigma\in\Sigma_{i_{0}}}(E_{i_{0}i_{0}}^{r(\sigma)}R)(\sigma^{-1})$. The following assertions hold:
@@
When $\Gamma$ is a group, Theorem 5.45 can be stated as follows: Let $G$ be a group, $D,D^{\prime}$ be $G$-graded division rings and $\overline{\sigma}:=(\sigma_{1},...,\sigma_{n})\in G^{n}$, $\overline{\delta}:=(\delta_{1},...,\delta_{m})\in G^{m}$. Then $\operatorname{M}_{n}(D)(\overline{\sigma})\cong_{gr}\operatorname{M}_{m}(D^{\prime})(\overline{\delta})$ if and only if $n=m$ and there exist a permutation $\pi$ of $\{1,...,n\}$ and $\tau\in G$ such that $D^{\prime}\cong_{gr}\operatorname{M}_{1}(D)(\tau^{-1})$ and $\delta_{\pi(i)}\in\tau(\operatorname{supp}D)\sigma_{i}$ for each i=1,,ni=1,...,n. This is essentially what was achieved in [14, p. 32-33]. Therefore, Theorem 5.45 generalizes the group graded case and we believe that our proof here, when adapted to $\Gamma$ being a group, provides a more elementary proof (or at least based on more elementary facts) of the result of [14]. ∎
@@ 6. The graded Jacobson-Chevalley density theorem
We begin by extending the definitions of gr-simple rings and modules for not necessarily object unital rings. Let $R$ be a $\Gamma$-graded ring and $M$ be a $\Gamma$-graded $R$-module. We say that $M$ is gr-simple if $MR\neq 0$ and the only graded submodules of $M$ are $\{0\}$ and $M$. And $M$ is said to be faithful if its right annihilator is zero, that is,
@@
Let $D$ be a $\Gamma$-graded division ring and $V$ be a $\Gamma$-graded unital left $D$-module. A graded subring $T$ of $\operatorname{END}_{D}(V)$ is said to be a gr-dense subring of $\operatorname{END}_{D}(V)$ if for all $n>0$, $\gamma_{1},...,\gamma_{n}\in\Gamma$, pseudo-linearly independent sequence $(u_{1},...,u_{n})\in V_{\gamma_{1}}\times\cdots\times V_{\gamma_{n}}$ and $(v_{1},...,v_{n})\in V_{\delta_{1}}\times\cdots\times V_{\delta_{n}}$ with $r(\delta_{i})=r(\gamma_{i})$ for all i=1,,ni=1,...,n there exists $t\in T$ such that $(u_{i})t=v_{i}$ for all i=1,,ni=1,...,n. By Corollary 4.5, $\operatorname{END}_{D}(V)$ is a gr-dense subring of $\operatorname{END}_{D}(V)$. In fact, if $(u_{1},...,u_{n})\in V_{\gamma_{1}}\times\cdots\times V_{\gamma_{n}}$ is a pseudo-linearly independent sequence and $(v_{1},...,v_{n})\in V_{\delta_{1}}\times\cdots\times V_{\delta_{n}}$ with $r(\delta_{i})=r(\gamma_{i})$ for all i=1,,ni=1,...,n, then, for each i=1,,ni=1,...,n, there exists $t_{i}\in\operatorname{END}_{D}(V)_{\gamma_{i}^{-1}\delta_{i}}$ such that $(u_{i})t_{i}=v_{i}$ and $(u_{j})t_{i}=0$ for every $j\neq i$. It suffices to take $t:=t_{1}+\cdots+t_{n}$.
@@
Now suppose that $n\geq 1$ and that the result holds for n1n-1. Let $_{D}V$ be a graded $D$-submodule of $_{D}S$ with $\operatorname{pdim}_{D}V=n$ and $x\in\operatorname{h}(S)\setminus V$. Let $(v_{1},...,v_{n})$ be a pseudo-basis of $_{D}V$. If $n>1$, define
@@
for all $a\in A$. Then the following homomorphism of $R$-modules is well-defined
@@
$(4)\implies(1):$ This implication follows from Theorem 5.23. ∎
@@ 7. Pseudo free module rings
    Conversely, suppose that $R$ is a gr-semisimple ring. Then $R$ is a pfm ring if and only if every gr-simple right $R$-module is pseudo-free.
@@
$\left\{\begin{array}[]{c}\textrm{graded division rings}\end{array}\right\}\subset\left\{\begin{array}[]{c}\textrm{pfm rings}\end{array}\right\}\subset\left\{\begin{array}[]{c}\textrm{gr-semisimple rings}\end{array}\right\}.$
@@
We point out that, in Example 7.2(2), $\mathcal{B}_{1}=\{\mathbb{I}_{(1,1)}\}$ and $\mathcal{B}_{2}=\{E_{13},\,E_{23}\}$ are two pseudo-bases of the $\Gamma$-graded $R$-module $S=R((1,1))$. Thus, there is no uniqueness of the cardinality of pseudo-bases of $\Gamma$-graded $R$-modules over pfm rings, and as a consequence, over gr-semisimple rings. Such uniqueness will characterize gr-division rings, but we will also be able to define an invariant similar to dimension for pfm rings, see Section 7.2.
@@ 7.1. Characterization of pfm rings
Fix $j\in J$ and $k\in K_{j}$. Set $R_{j}:=\operatorname{M}_{K_{j}}(D_{j})(\overline{\sigma}_{j})$ and let $\iota_{j}:R_{j}\to R$ be the canonical inclusion. Since the $\Gamma$-graded right $R$-module $\iota_{j}(E_{kk}^{e_{j}}R_{j})$ has a pseudo-basis, there exist $\gamma_{j}\in\Gamma$ and $a_{j}\in(E_{kk}^{e_{j}}R_{j})_{\gamma_{j}}$ such that $\iota_{j}(a_{j})\cdot r\neq 0$ for all $r\in 1_{d(\gamma_{j})}R\setminus\{0\}$. In particular, $a_{j}\neq 0$ and, thus, $K_{j,d(\gamma_{j})}\neq\emptyset$. Let $j^{\prime}\in J$ be such that there exists $k^{\prime}\in K_{j^{\prime},d(\gamma_{j})}$. Then $0\neq E_{k^{\prime}k^{\prime}}^{e_{j^{\prime}}}\in(R_{j^{\prime}})_{d(\gamma_{j})}$ and it follows that $\iota_{j}(a_{j})\iota_{j^{\prime}}(E_{k^{\prime}k^{\prime}}^{e_{j^{\prime}}})\neq 0$. But this is only possible if $j=j^{\prime}$. Hence, $\{j^{\prime}\in J:K_{j^{\prime},f_{j}}\neq\emptyset\}=\{j\}$. Now observe that $a_{j}\cdot r_{j}\neq 0$ for all $r_{j}\in R_{j}(d(\gamma_{j}))\setminus\{0\}$. Proceeding as in the proof of $(2)\implies(3)$ of Theorem 7.3, we get that $K_{j,d(\gamma_{j})}$ contains exactly one element.
@@
$(6)\implies(2)$: Let $j\in J$. By Theorem 5.23, there exist a non-empty set $K_{j}$, $e_{j}\in\Gamma_{0}$, a $\Gamma$-graded division ring $D_{j}$ with $\operatorname{supp}(D_{j})\subseteq e_{j}\Gamma e_{j}$ and a $d$-finite sequence $\overline{\sigma}_{j}:=(\sigma_{jk})_{k\in K_{j}}\in(e_{j}\Gamma)^{K_{j}}$ such that
@@ 7.2. Invariance of the number of elements of pseudo-bases
As we are going to see next, it is still possible to define an invariant for graded modules over pfm rings similar to the pseudo-dimension and that coincides with the graded length of a finitely generated graded module over a pfm ring. To that end, we begin by pointing out some facts about gr-semisimple modules.
@@
Suppose now that $X$ is a $\Gamma$-graded module. We say that a pseudo-linearly independent sequence $(x_{i})_{i\in I}$ of homogeneous elements of $X$ is a gr-simple sequence if $x_{i}R$ is a gr-simple $R$-module for all $i\in I$. If, moreover, $(x_{i})_{i\in I}$ is a pseudo-basis of $X$ we say that it is a gr-simple pseudo-basis of $X$. In this event, we will write $\operatorname{spdim}_{R}(X)=|I|$. We proceed to show that $\operatorname{spdim}_{R}(X)$ is a well-behaved invariant for pfm rings.
@@
Let $R=\bigoplus\limits_{\gamma\in\Gamma}R_{\gamma}$ be a pfm ring and $M=\bigoplus\limits_{\gamma\in\Gamma}M_{\gamma}$ be a $\Gamma$-graded $R$-module. The following assertions hold:
@@
    If $N$ is a graded submodule of $M$, then spdim_R(N)+spdim_R(M/N)=spdim_R(M).
@@ 7.3. More on gr-division rings
    There exist a set $J$, a family of non-empty subsets $\{K_{j}:j\in J\}$, a sequence of idempotents $(e_{j})_{j\in J}\in(\Gamma_{0})^{J}$ such that
@@
$(3)\implies(4)$: Straightforward.
$(4)\implies(5)$: By Theorem 5.30, there exist families $\{K_{j}:j\in J\}$ and $\{e_{j}:j\in J\}$ of non-empty sets and idempotents of $\Gamma$, respectively, and there exist, for each $j\in J$, a $\Gamma$-graded division ring $D_{j}$ with $\operatorname{supp}(D_{j})\subseteq e_{j}\Gamma e_{j}$ and a $d$-finite sequence $\overline{\sigma}_{j}:=(\sigma_{jk})_{k\in K_{j}}\in(e_{j}\Gamma)^{K_{j}}$, such that the family $\{\operatorname{M}_{K_{j}}(D_{j})(\overline{\sigma}_{j}):j\in J\}$ is summable and
@@
$(5)\implies(6)$: For each $j\in J$, $R_{j}:=\operatorname{M}_{K_{j}}(D_{j})(\overline{\sigma}_{j})$ is a gr-prime gr-division ring, by Theorem 4.2. Suppose now that $j,j^{\prime}\in J$ and $\gamma\in\operatorname{supp}(R_{j})\cap\operatorname{supp}(R_{j^{\prime}})$. Since $|K_{j,e}|=1$ for all $e\in\Gamma^{\prime}_{0}(R)$, there exists a unique $p\in K_{j}$ such that $d(\sigma_{jp})=r(\gamma)$. Analogously, there exists a unique $p^{\prime}\in K_{j^{\prime}}$ such that $d(\sigma_{j^{\prime}p^{\prime}})=r(\gamma)$. Thus, we obtain non-empty $K_{j,r(\gamma)}$ and $K_{j^{\prime},r(\gamma)}$, hence $j,j^{\prime}\in J_{r(\gamma)}$ and, therefore, $j^{\prime}=j$.
@@
The next result shows that to be a pfm ring is equivalent to be a gr-division ring for object crossed products.
@@
The following assertions hold:
@@ 8. Semisimple categories
Many of the results in this section can be proved in a more general context. The authors have been working on a paper in which they introduce the concept of groupoid graded categories, which generalize group graded categories, and results are obtained for such more general categories.
@@
In [32], $\mathcal{C}(A,-)$ and $\mathcal{C}(-,A)$ are denoted by $\mathcal{C}_{A}$ and $\mathcal{C}_{A}^{~*}$, respectively. In [32, p. 18], a right (left) ideal of $\mathcal{C}$ is defined as a subfunctor of $\mathcal{C}_{A}$ (of $\mathcal{C}_{A}^{~*}$) for some $A\in\mathcal{C}_{0}$. We point out that we are using the opposite notation because our composition of morphisms in the category $\mathcal{C}$ is the usual from right to left, while [32, p. 7] composes morphisms of $\mathcal{C}$ from left to right.
In what follows, inspired by [18, Proposition 2 (p. 347)], we present a possible motivation for the previous definitions.
@@
If $M=M(\varepsilon_{A})$ for some $A\in\mathcal{C}_{0}$, then we will write M(,A)M(-,A) instead of $M(-,\mathcal{C}_{0})$.
@@
The next result links the constructions of the previous paragraphs. Recall that, by Remark 2.5(1), every unital right $R[\mathcal{C}]$-module $M$ can be regarded as a $\mathcal{G}$-graded $R[\mathcal{C}]$-module making $M=M(\varepsilon_{A})$ for some $\varepsilon_{A}\in\mathcal{G}_{0}$. Therefore, Theorem 8.1(1) is a rephrasing of [18, Proposition 2 (p. 347)].
@@
Let $\mathcal{G}:=\mathcal{C}_{0}\times\mathcal{C}_{0}$. The following functors define an equivalence of categories
$\mathcal{G}-\operatorname{gr\textrm{-}\!}R[\mathcal{C}]$ $\leftrightarrows\operatorname{Fun_{gr}}(\mathcal{C}^{op},\mathcal{G}\varepsilon-\mathbb{Z}^{(\mathcal{G}_{0})}\operatorname{\!\textrm{-}mod})$
$M$ $\mapsto M(-,\mathcal{C}_{0})$
$M[F]$ $\mapsfrom F.$
Moreover, if $A\in\mathcal{C}_{0}$, then such equivalence induces
@@
    an equivalence of categories
    $\varepsilon_{A}\mathcal{G}-\operatorname{gr\textrm{-}\!}R[\mathcal{C}]$ $\leftrightarrows\operatorname{Fun}(\mathcal{C}^{op},\mathcal{A}b)$
    $M$ $\mapsto M(-,A)$
    $M[F]$ $\mapsfrom F;$
@@
    sending $R[\mathcal{C}](\varepsilon_{A})$ to $\mathcal{C}(-,A)$.
@@
${MI_{X}=M(X,\mathcal{C}_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{X}}$${M(Y,\mathcal{C}_{0})=MI_{Y}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{M(f,\mathcal{C}_{0})}$$\scriptstyle{\alpha_{Y}}$${NI_{X}=N(X,\mathcal{C}_{0})}$${N(Y,\mathcal{C}_{0})=NI_{Y}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{N(f,\mathcal{C}_{0})}$
@@
It is straightforward to check that $M[M(-,\mathcal{C}_{0})]\cong_{gr}M$ for each object $M$ of $\mathcal{G}-\operatorname{gr\textrm{-}\!}R[\mathcal{C}]$, and $M[F](-,\mathcal{C}_{0})\cong F$ for each object $F$ of $\operatorname{Fun_{gr}}(\mathcal{C}^{op},\mathcal{G}\varepsilon-\mathbb{Z}^{(\mathcal{G}_{0})}\operatorname{\!\textrm{-}mod})$. Furthermore, such isomorphisms are natural.
(1) Fix $A\in\mathcal{C}_{0}.$ For $F\in\operatorname{Fun}(\mathcal{C}^{op},\mathcal{A}b)$, consider $M[F]$ as an object of $\varepsilon_{A}\mathcal{G}-\operatorname{gr\textrm{-}\!}R[\mathcal{C}]$ via $M[F]_{(A,X)}:=F(X)$ for each $X\in\mathcal{C}_{0}$. Then the assignments $M\mapsto M(-,A)$ and $F\mapsto M[F]$ define an equivalence between the categories $\varepsilon_{A}\mathcal{G}-\operatorname{gr\textrm{-}\!}R[\mathcal{C}]$ and $\operatorname{Fun}(\mathcal{C}^{op},\mathcal{A}b)$.
(2) Note that if $M$ is a graded right ideal of $R[\mathcal{C}]$ with $M=M(\varepsilon_{A})$ for some $A\in\mathcal{C}_{0}$, then M(,A)M(-,A) is a subfunctor of $\mathcal{C}(-,A)$, i.e., a right sieve on $A$ in $\mathcal{C}$. And, if $F$ is a right sieve on $A$ in $\mathcal{C}$, then $M[F]$ is naturally gr-isomorphic to a graded right ideal of $R[\mathcal{C}]$ contained in $R[\mathcal{C}](\varepsilon_{A})$. Thus, (2) follows from (1). ∎